Circular No. 7275

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ef IN UGC 607 Jean Mueller reports her discovery of a supernova (mag ∼ 17.5) on a red

plate taken by K. Rykoski and herself on Oct. 9 UT with the 1.2-m Oschin Schmidt Telescope in the course of the Palomar Outer Solar System Ecliptic (POSSE) Survey. SN 1999ef is located at α = 0^{h}58^{m}46^{s}.2, δ = +12^{o}44′45′′ (equinox 2000.0), which is 20′′ east and 10′′ south of the nucleus of UGC 607. There is no object at this position on the Digitized Sky Survey or on second Palomar Sky Survey plates. SN 1999ef also appears on a POSSE plate taken on Oct. 10.

M. J. Kuchner, Astronomy Department, California Institute of Technology, reports that a spectrum was taken by A. Bouchez and himself on Oct. 10 UT with the 5-m Hale telescope (+ double spectrograph) at Palomar. D. Branch identifies the object as a normal post-maximum type-Ia supernova from the spectrum.

SUPERNOVA 1999eg IN IC 1861 G. M. Hurst, Basingstoke, England, reports the discovery by Mark

Armstrong, Rolvenden, of an apparent supernova (mag ∼ 17.5) on an unfiltered CCD frame taken on Oct. 4.078 UT with a 0.30-m reflector in the course of the U.K. Nova/Supernova Patrol. In response to an alert, N. James, Chelmsford, recorded a very weak image on Oct. 10.87 (0.3-m reflector). Further images by S. Foulkes, Ashperton Village, and Armstrong on Oct. 11 confirmed the presence of the object. Armstrong measured the following precise position from his latest image on Oct. 11.875: α = 2^{h}53^{m}08^{s}.39, δ = +25^{o}29′24′′.2 (equinox 2000.0); the discovery image yields position end figures 08^{s}.45, 23′′.7. Hurst notes that the object is absent from the Palomar Sky Surveys, and no candidate appears in the USNO A2.0 catalogue.

SUPERNOVAE 1999by, 1999dk, 1999dq Unfiltered CCD magnitudes from L. Kiss, K. Sárneczky, and G. Szabó,

Piszkéstető, Konkoly Observatory: SN 1999by in NGC 2841, Sept. 24.07 UT, 18.6 ± 0.3; SN 1999dk in UGC 1087, 23.88, 15.7 ± 0.3; SN 1999dq in NGC 976, 23.87, 14.7 ± 0.2.

1999 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green